renee is 3 times as old as dana. five years from now, renee will be twice as old as dana. how old are they now?
Answer by mananth(16946)

Let dana'a age be x
renee's age =3x
five years later they will be
x+5 & 3x + 5
3x+5= 2*(x+5)
3x+5=2x+10
3x-2x=10-5
x=5 years -dana's age
renee's age = 3x = 15 years